OpenStudy (anonymous):

The length of a rectangle is 4yd longer than its width. If the perimeter of the rectangle is 48yd, find its area.

OpenStudy (nurali):

the P of a rectangle is the sum of all 4 sides or is generaly written as: 2*l + 2*w = P and area is wrtten as l*w. . The question states the following to be true: Let w represent the width l = (w + 4) . 2*(w + 4) + 2*w = 40yds . Simplify and solve for w 4w + 8 = 48yds 4w = 40yds w= 10 yds l = 14yds . A = l*w A=14*10 A=140 yd^2